Find the 6th term of geometric sequence with the given terms
A4=-12 a5=-4

Answer by richwmiller(17219)

Tn = t1 * r^(n - 1)
-4 = t1 * r^(4)
-12= t1 * r^(3)
r=1/3
t6 = t1 * r^(5)
r = 1/3, t1 = -324, t6 = -4/3